find equation of the line which is perpendicular to the line y=5x+9 and has y - intercept=6
Answer by mananth(16946)

1 y = 5 x + 9
Divide by 1
y = 5 x + 9
Compare this equation with y=mx+b
slope m = 5.00

The slope of a line perpendicular to the above line will be the negative reciprocal
m1*m2=-1
m= - 1/5 ,point ( 0 , 6 )
y intercept =6
The required equation isy =- 1/ 5 x + 6
m.ananth@hotmail.ca